# SWAT: Sliding Window Adversarial Training for Gradual Domain Adaptation

### v1

@@ 4.4 The Overall Objective for Gradual Domain Adaptation
where $l$ is the cross-entropy loss. The difference with previous GDA methods is that when $h$ comes from the unlabeled domain, $\hat{y}_{t}(x)$ is the pseudo-label generated by the model $g$. But when $h$ is a feature generated by $G_{m}(h_{0})$, it represents the ground-truth label of the original sample $h_{0}$ from the source domain.
The overall objective of the sliding window adversarial training is formulated as follows:
@@
where $\mathcal{L}_{l}$ is the adversarial training loss which defines as: $\mathcal{L}^{l}=\mathcal{L}^{l}_{adv}+\mathcal{L}^{l}_{cycle}+\mathcal{L}^{l}_{st}$. By optimizing Eq. 8, we achieve continuous flow matching in the feature space. To best understand our method, we summarize the pseudo-code of SWAT in Algorithm 1 and provide an illustration of its architecture in Fig. 5.
@@
[…]
@@
8:     for $p=0$ to $1$, with step size $\Delta p$ do
@@
10:        $G_{m},G_{s}\leftarrow\text{Adam}(\theta_{G_{m}},\theta_{G_{s}})$
11:        $D_{s},D_{l},D_{r}\leftarrow\text{Adam}(\theta_{D_{s}},\theta_{D_{l}},\theta_{D_{r}})$
12:        $f,g\leftarrow\text{Adam}(\theta_{f},\theta_{g})$
13:     end for
14:  end for
@@ 5 Experiments
5.1 Datasets
@@
Constructed from MNIST (Deng, 2012), this dataset (He et al., 2023) contains 50,000 source domain images (original digits) and 50,000 target domain images rotated by 45°. Intermediate domains interpolate rotation angles between 0° and 45°.
@@
MNIST images are normalized to [0,1] for the source domain and shifted to [1,2] for the target domain (He et al., 2023), with intermediate domains generated by linearly interpolating color intensity.
Portraits (Ginosar et al., 2015)
Chronologically divided into 9 temporal domains (1905–2013), each with 2,000 images (Kumar et al., 2020). The first and last domains serve as source/target; images are resized to 32×32 pixels.
Cover Type(Blackard, 1998)
Sorted by horizontal distance to water, this tabular dataset (Kumar et al., 2020) uses 50,000 source samples, 10×40,000 intermediate domains, and 50,000 target samples for classifying spruce fir vs. Rocky Mountain pine.
5.2 Implementation Details
Table 1: Comparison of GDA methods and UDA methods on datasets.
UDA/GDA methods Rotated MNIST Color-Shift MNIST Portraits Cover Type
DANN (Ganin et al., 2016) $44.23$ $56.5$ $73.8$
DeepCoral (Sun & Saenko, 2016) $49.6$ $63.5$ $71.9$
DeepJDOT (Damodaran et al., 2018) $51.6$ $65.8$ $72.5$
GST (Kumar et al., 2020) $83.8$ $74.0$ $82.6$ $73.5$
CNF (Sagawa & Hino, 2025) $62.55$ $84.57$
IDOL (Chen & Chao, 2021) $87.5$ $85.5$
GGF (Zhuang et al., 2024) $67.72$ $86.16$
GOAT (He et al., 2023) $86.4$ $91.8$ $83.6$ $69.9$
SWAT (Ours) $\mathbf{96.7}$ $\mathbf{99.6}$ $\mathbf{87.4}$ $\mathbf{75.0}$
We design distinct architectures for image and tabular data. For Rotated MNIST, Color-Shift MNIST, and Portraits, a CNN with three 32-channel convolutional layers is used, followed by a classifier with two 256-unit hidden layers. For Cover Type, a fully connected network with increasing hidden dimensions (128→256→512) and ReLU activations is adopted.
The transport module employs generators with residual blocks (three linear layers) and discriminators with three 128-unit linear layers. Optimization uses Adam (Kingma & Ba, 2014), with Dropout (Srivastava et al., 2014) and Batch Normalization (Ioffe & Szegedy, 2015) for regularization. The number of intermediate domains (0–4) is treated as a hyperparameter. All experiments run on NVIDIA RTX 4090 GPUs. Following (Kumar et al., 2020), we filter 10% of low-confidence data by computing the margin between the top two predicted probabilities.
Compared with the text on UDA methods by He et al. (2023), GDA methods are superior as shown in Table 1. For more detailed experimental steps, please refer to Appendix A.
5.3 Results
Table 2: Comparison of the accuracy of our method with the state-of-the-art methods GST(Kumar et al., 2020), GOAT(He et al., 2023) on vision datasets (Rotated MNIST, Color-Shift MNIST, and Portraits datasets), along with 68% confidence intervals of the mean over 5 runs.
# Given Rotated MNIST Color-Shift MNIST Portraits
Domains GST GOAT SWAT GST GOAT SWAT GST GOAT SWAT
2 $54.9\pm 0.2$ $53.5\pm 1.0$ $\mathbf{88.1\pm 1.5}$ $27.0\pm 0.3$ $72.0\pm 6.0$ $\mathbf{98.8\pm 0.3}$ $75.0\pm 1.7$ $78.6\pm 2.2$ $\mathbf{85.3\pm 0.1}$
3 $60.0\pm 0.3$ $57.2\pm 0.3$ $\mathbf{96.1\pm 0.1}$ $34.2\pm 1.7$ $83.4\pm 2.9$ $\mathbf{99.5\pm 0.0}$ $75.1\pm 1.0$ $80.2\pm 1.3$ $\mathbf{84.8\pm 1.0}$
4 $67.2\pm 0.6$ $68.4\pm 1.4$ $\mathbf{96.4\pm 0.0}$ $55.0\pm 1.9$ $89.1\pm 3.6$ $\mathbf{99.6\pm 0.0}$ $78.4\pm 0.9$ $80.5\pm 1.3$ $\mathbf{86.1\pm 0.3}$
5 $71.9\pm 0.8$ $78.8\pm 0.8$ $\mathbf{96.5\pm 0.2}$ $66.8\pm 2.2$ $94.9\pm 1.0$ $\mathbf{99.6\pm 0.0}$ $76.4\pm 1.8$ $79.4\pm 0.6$ $\mathbf{87.0\pm 0.0}$
6 $75.6\pm 1.4$ $85.8\pm 0.9$ $\mathbf{96.7\pm 0.1}$ $74.0\pm 3.4$ $95.7\pm 0.3$ $\mathbf{99.6\pm 0.0}$ $80.9\pm 0.6$ $83.1\pm 0.6$ $\mathbf{87.4\pm 0.2}$
Table 3: Comparison of the accuracy of our method with the state-of-the-art methods GST(Kumar et al., 2020), GOAT(He et al., 2023) on Tabular datasets (Cover Type datasets), along with 68% confidence intervals of the mean over 5 runs.
# Given Cover Type
Domains GST GOAT SWAT
2 $69.1\pm 0.1$ $69.0\pm 0.0$ $\mathbf{75.0\pm 0.0}$
3 $71.1\pm 0.2$ $69.0\pm 0.0$ $\mathbf{74.3\pm 0.2}$
4 $72.4\pm 0.1$ $69.0\pm 0.0$ $\mathbf{74.6\pm 0.1}$
5 $72.8\pm 0.1$ $69.1\pm 0.1$ $\mathbf{74.6\pm 0.1}$
6 $\mathbf{73.1\pm 0.1}$ $69.3\pm 0.0$ $\mathbf{73.7\pm 0.2}$
Table 2 and Table 3 presents the comparison of our method, SWAT, with GST (Kumar et al., 2020) and GOAT (He et al., 2023) across vision and tabular datasets, including Rotated MNIST, Color-Shift MNIST, Portraits, and Cover Type. The table presents the test results for the same dataset processed with the same encoder classifier and the same low confidence sample selection strategy. Each result is the average of five random seeds and the variance.
For the vision datasets (Rotated MNIST, Color-Shift MNIST, and Portraits), SWAT outperforms GST and GOAT across all the given domain settings (2, 3, 4, 5, and 6 domains). The performance improvement is particularly noticeable in scenarios where the number of given domains is small. In the case of the Cover Type dataset, SWAT also demonstrates significant improvements. The confidence intervals in the results are relatively narrow, indicating that the reported accuracies are consistent across multiple runs.
By leveraging domain flow and feature transfer more efficiently than both GST and GOAT, SWAT is able to adapt better across diverse datasets. In particular, the enhancement in performance with fewer domains (such as 2 or 3 domains) suggests that SWAT may benefit from more efficient utilization of available domain-specific data, which is crucial when only a limited number of domains are available for training.
Overall, SWAT consistently outperforms both GST and GOAT across all datasets, demonstrating its effectiveness in transferring knowledge between domains and improving model generalization. These results highlight the advantages of our proposed approach, particularly in datasets where domain flow and feature transfer are critical for achieving high accuracy. For more detailed experimental results, please refer to Appendix A.
5.4 Inter-Domain Distance Reduction Analysis
Our experimental results quantitatively and visually validate that the proposed Sliding Window Adversarial Training (SWAT) effectively reduces domain gaps while preserving semantic consistency. Fig. LABEL:fig:feature_encoder illustrates this through two complementary perspectives.
Quantitative Analysis via $\mathcal{A}$-Distance:
As shown in the left panel of Fig. LABEL:fig:feature_encoder, the $\mathcal{A}$-distance (Ben-David et al., 2010) between the source domain $\mathcal{H}_{0}$ (red curve) and target domain $\mathcal{H}_{n}$ (blue curve) exhibits drastic fluctuations (peak at 1.498) when using a fixed encoder. This indicates unstable feature alignment under large domain shifts. In contrast, our SWAT framework (orange curve) maintains near-zero distances ($<0.11$) to critical intermediate domains $\mathcal{H}_{l},\mathcal{H}_{r}$ across all positions, achieving a 63.7% reduction in the average $\mathcal{A}$-distance between $\mathcal{H}_{0}$ and $\mathcal{H}_{n}$ (0.104 vs. 1.284). The symmetrical decay of bidirectional distances confirms balanced adaptation between forward and backward domain transitions. Orange curve achieves near-zero distances to adjacent critical domains $\mathcal{H}_{l},\mathcal{H}_{r}$ throughout the adaptation path, demonstrating smooth domain transitions.
Feature Visualization Analysis
The T-SNE visualizations (Fig. LABEL:fig:feature_encoder) quantitatively reveal the geometric impact of adaptation strategies:
1. Unconstrained Alignment Failure: Direct mapping to $\mathcal{H}_{n}$ sans flow matching (Fig. LABEL:fig:feature_encoder:b) causes catastrophic cluster overlap, increasing class impurity by 38.7% compared to $\mathcal{H}_{0}$, as rigid alignment disrupts local semantic structures.
2. Geodesic Flow Preservation: Our flow-constrained adaptation (Fig. LABEL:fig:feature_encoder:c) maintains 89.4% of $\mathcal{H}_{0}$’s cluster purity (adjusted Rand index 0.82 vs. baseline 0.41) through adversarial flow matching that preserve isometric relationships between neighboring domains $\mathcal{H}_{l}\leftrightarrow\mathcal{H}_{r}$.
3. Adversarial Trajectory Smoothing: The non-adversarial path $\mathcal{H}_{0}\rightarrow\mathcal{H}_{n}$ (Fig. LABEL:fig:feature_encoder:d) exhibits discontinuous jumps (Hausdorff distance 4.72), while our adversarial flow $\mathcal{H}_{0}\rightarrow\mathcal{H}_{g}\rightarrow\mathcal{H}_{n}$ (Fig. LABEL:fig:feature_encoder:e) reduces trajectory fragmentation by 75.6% (Hausdorff 1.15).
This geometric perspective explains SWAT’s dual advantage: Semantic invariance via optimal transport between adjacent domains; Topological continuity through adversarial curvature control in feature space.
@@ 5.4 Inter-Domain Distance Reduction Analysis
000.50.50.50.51111000.050.050.050.050.10.10.10.1
Figure 6: The figure illustrates the $\mathcal{A}$-distance (Ben-David et al., 2010; Mansour et al., 2009) to the source domain (the red line) and the target domain (the blue line) for various domains. The orange line represents the distance between the left and right critical domains and the intermediate domains after cross-domain transport. On the left, the $\mathcal{A}$-distance is computed with a fixed encoder in a self-training setup, while on the right, the distance is calculated using our SWAT representation with a flow encoder. (Explanation: Four given fields with two-step iteration in SWAT.)

### v2

@@ 4.1 The Overall Objective
Figure 3: $\mathcal{H}^{(l+p)}$ lies along the shortest domain flow determined by $p$.
where $l$ is the cross-entropy loss. When $h$ comes from the unlabeled domain, $\hat{y}_{t}(x)$ is the pseudo-label generated by the model $g$. But when $h$ is a feature generated by $G_{m}(h_{0})$, it represents the ground-truth label of the original representation $h_{0}$ from the source domain.
As shown in Fig. 3, the generated feature space $\mathcal{H}^{(l+p)}$ is enforced to satisfy the condition $\text{dist}(\mathcal{H}_{l},\mathcal{H}^{(l+p)})$$\text{dist}(\mathcal{H}_{r},\mathcal{H}^{(l+p)})$==p/(1p)p/(1-p), where $\text{dist}(\cdot,\cdot)$ denotes a valid distance metric between two distributions.
The overall objective of SWAT is formulated as follows:
@@
where $\mathcal{L}_{l}$ is the adversarial training loss which defines as: $\mathcal{L}^{l}=\mathcal{L}^{l}_{adv}+\mathcal{L}^{l}_{cycle}+\mathcal{L}^{l}_{st}$. By optimizing Eq. 13, we achieve continuous flow matching in the feature space. For clear understanding, we summarize the main idea of SWAT in Algorithm 1.
@@
1:Input: A series of domains $\mathcal{D}_{0}\text{(source)},$ $\mathcal{D}_{1},$ $\ldots,$ $\mathcal{D}_{n-1}$ $\mathcal{D}_{n}\text{(target)}$, pretrained encoder $f$ and classifier $g$.
@@
8:  for $p=0$ to 11, with step size $\Delta p$ do
@@
10:   update $G_{m},G_{s},D_{s},D_{l},D_{r},f,g$
11:  end for
12:end for
@@ 5 Experiments
5.1 Datasets and Implementation Details
Following the standard GDA protocol, we conduct extensive experiments on 6 datasets. Rotated MNIST is constructed from MNIST [27], this dataset contains 50,000 source domain images (original digits) and 50,000 target domain images rotated by 45°. Intermediate domains interpolate rotation angles between 0° and 45°. Color-Shift MNIST images are normalized to [0,1] for the source domain and shifted to [1,2] for the target domain [13], with intermediate domains generated by linearly interpolating color intensity. Portraits [28] are chronologically divided into 9 temporal domains (1905–2013), each with 2,000 images [7]. The first and last domains serve as source/target; images are resized to 32×32 pixels. Cover Type[29] tabular dataset sorted by horizontal distance to water, uses 50,000 source samples, 10×40,000 intermediate domains, and 50,000 target samples [7] for classifying spruce fir vs. Rocky Mountain pine.
To evaluate the performance of GDA methods under high-severity shifts, we introduce a new evaluation protocol using the corruption benchmarks CIFAR-10C, CIFAR-100C [30]. Each benchmark applies 15 corruption types at five severity levels to the validation and test splits of CIFAR [31]. We regard the clean training images as the source domain, the images of severity levels 1–4 (across all corruption types) as a sequence of intermediate domains, and treat severity level 5 as the target domain. Following the RobustBench benchmark [32, 33], WideResNet-28 [34] and ResNeXt-29 [35] used as the source model for CIFAR10-to-CIFAR10C and CIFAR100-to-CIFAR100C, respectively.
All results are averaged over 5 runs. Please refer to Appendix A.2 for more detailed implementation.
5.2 Experimental Results
Table 1: Comparison of domain adaptation methods on 4 GDA datasets.
Methods Gradual Rotated MNIST Color-Shift MNIST Portraits Cover Type
DANN [36] $44.2$ $56.5$ $73.8$ -
DeepCoral [37] $49.6$ $63.5$ $71.9$ -
DeepJDOT [38] $51.6$ $65.8$ $72.5$ -
GST [7] (ICML’20) $83.8$ $74.0$ $82.6$ $73.5$
IDOL [39] (NeurIPS’21) $87.5$ - $85.5$ -
AGST [40] (IEEE’22) 76.2 - 77.6
GGF [22] (ICLR’24) $67.7$ - $86.2$ -
GOAT [13] (JMLR’24) $86.4$ $91.8$ $83.6$ $69.9$
DRO [41] (NeurIPS’24) $53.2$ - - -
AST [10] (NeurIPS’24) $90.6$ - $84.8$ -
CNF [42] (Neural Computation’25) $62.6$ - $84.6$ -
SWAT (Ours) $\mathbf{96.7}$ $\mathbf{99.6}$ $\mathbf{87.4}$ $\mathbf{75.0}$
Compared with the UDA methods [13], table 1 highlights the clear advantage of gradual domain adaptation (GDA) over traditional unsupervised domain adaptation (UDA) and reveals the clear performance advantages in the GDA setting: Whereas UDA methods like DANN, DeepCORAL, and DeepJDOT struggle under large domain shifts, achieving at best 51.6% on Rotated MNIST and 65.8% on Color-Shift MNIST, these results highlight SWAT’s ability to learn smooth and discriminative representations under gradual shifts.
Table 2: Comparison of SWAT and other GDA methods on 4 datasets given different numbers of intermediate domains.
Given Domains
@@ 5.2 Experimental Results
Given Domains
@@
GST GOAT SWAT GST GOAT SWAT
2 $54.9\pm 0.2$ $53.5\pm 1.0$ $\mathbf{88.1\pm 1.5}$ 2 $27.0\pm 0.3$ $72.0\pm 6.0$ $\mathbf{98.8\pm 0.3}$
3 $60.0\pm 0.3$ $57.2\pm 0.3$ $\mathbf{96.1\pm 0.1}$ 3 $34.2\pm 1.7$ $83.4\pm 2.9$ $\mathbf{99.5\pm 0.0}$
4 $67.2\pm 0.6$ $68.4\pm 1.4$ $\mathbf{96.4\pm 0.0}$ 4 $55.0\pm 1.9$ $89.1\pm 3.6$ $\mathbf{99.6\pm 0.0}$
@@
[…]
Given Domains
Portraits
Given Domains
Cover Type
GST GOAT SWAT GST GOAT SWAT
2 $75.0\pm 1.7$ $78.6\pm 2.2$ $\mathbf{85.3\pm 0.1}$ 2 $69.1\pm 0.1$ $69.0\pm 0.0$ $\mathbf{75.0\pm 0.0}$
3 $75.1\pm 1.0$ $80.2\pm 1.3$ $\mathbf{84.8\pm 1.0}$ 3 $71.1\pm 0.2$ $69.0\pm 0.0$ $\mathbf{74.3\pm 0.2}$
4 $78.4\pm 0.9$ $80.5\pm 1.3$ $\mathbf{86.1\pm 0.3}$ 4 $72.4\pm 0.1$ $69.0\pm 0.0$ $\mathbf{74.6\pm 0.1}$
5 $76.4\pm 1.8$ $79.4\pm 0.6$ $\mathbf{87.0\pm 0.0}$ 5 $72.8\pm 0.1$ $69.1\pm 0.1$ $\mathbf{74.6\pm 0.1}$
6 $80.9\pm 0.6$ $83.1\pm 0.6$ $\mathbf{87.4\pm 0.2}$ 6 $\mathbf{73.1\pm 0.1}$ $69.3\pm 0.0$ $\mathbf{73.7\pm 0.2}$
Table 2 compares SWAT against GST [7] and GOAT [13] on both vision benchmarks (Rotated MNIST, Color-Shift MNIST, Portraits) and the Cover Type tabular dataset, using the same encoder–classifier architecture and low-confidence sample selection strategy throughout. SWAT consistently outperforms GST and GOAT across every setting, with the largest gains observed when only two or three domains are available. Narrow confidence intervals further confirm the stability of our results. By more effectively leveraging domain flow and feature transfer SWAT delivers superior adaptation across diverse data modalities. (Additional experiments and computational-cost comparisons are provided in Appendices A.1 and C, respectively.)
Table 3: Comparison of classification error rates (%) at severity level 5 for test‐time adaptation (TTA) and gradual domain adaptation (GDA) methods on CIFAR-10C and CIFAR-100C. Lower is better.
Method
Gradual
gaussian
shot
impulse
defocus
glass
motion
zoom
snow
frost
fog
brightness
contrast
elastic
pixelate
jpeg
Mean
CIFAR-10C
Source only 72.3 65.7 72.9 46.9 54.3 34.8 42.0 25.1 41.3 26.0 9.3 46.7 26.6 58.5 30.3 43.5
BN-1 28.1 26.1 36.3 12.8 35.3 14.2 12.1 17.3 17.4 15.3 8.4 12.6 23.8 19.7 27.3 20.4
TENT-cont. [43] 24.8 20.6 28.6 14.4 31.1 16.5 14.1 19.1 18.6 18.6 12.2 20.3 25.7 20.8 24.9 20.7
AdaContrast [44] 29.1 22.5 30.0 14.0 32.7 14.1 12.0 16.6 14.9 14.4 8.1 10.0 21.9 17.7 20.0 18.5
CoTTA [45] 24.3 21.3 26.6 11.6 27.6 12.2 10.3 14.8 14.1 12.4 7.5 10.6 18.3 13.4 17.3 16.2
GTTA-MIX [46] 23.4 18.3 25.5 10.1 27.3 11.6 10.1 14.1 13.0 10.9 7.4 9.0 19.4 14.5 19.8 15.6
GST [7] 50.0 43.9 50.3 20.6 51.2 17.2 16.7 17.5 24.3 17.5 6.9 13.2 24.9 39.9 26.6 28.1
GOAT [13] 72.7 65.7 73.0 46.7 54.5 34.3 41.5 24.9 41.0 26.0 9.3 46.6 26.4 58.1 30.2 43.4
SWAT(ours) 21.4 20.0 26.8 9.7 28.5 10.2 8.4 3.1 13.4 11.5 7.1 8.9 19.8 13.3 20.1 15.4
CIFAR-100C
Source only 73.0 68.0 39.4 29.3 54.1 30.8 28.8 39.5 45.8 50.3 29.5 55.1 37.2 74.7 41.2 46.4
BN-1 42.1 40.7 42.7 27.6 41.9 29.7 27.9 34.9 35.0 41.5 26.5 30.3 35.7 32.9 41.2 35.4
TENT-cont. [43] 37.2 35.8 41.7 37.9 51.2 48.3 48.5 58.4 63.7 71.1 70.4 82.3 88.0 88.5 90.4 60.9
AdaContrast [44] 42.3 36.8 38.6 27.7 40.1 29.1 27.5 32.9 30.7 38.2 25.9 28.3 33.9 33.3 36.2 33.4
CoTTA [45] 40.1 37.7 39.7 26.9 38.0 27.9 26.4 32.8 31.8 40.3 24.7 26.9 32.5 28.3 33.5 32.5
GTTA-MIX [46] 36.4 32.1 34.0 24.4 35.2 25.9 23.9 28.9 27.5 30.9 22.6 23.4 29.4 25.5 33.3 28.9
GST [7] 49.8 56.7 32.3 22.5 41.6 25.0 23.3 30.3 32.2 38.1 22.1 27.0 33.1 40.8 35.8 33.3
GOAT [13] 73.4 67.9 39.1 28.7 53.8 30.2 28.7 39.3 45.7 50.0 29.4 53.7 36.8 74.3 41.2 46.2
SWAT(ours) 28.6 26.9 23.5 22.3 29.0 22.7 22.4 24.4 24.3 25.7 21.5 22.7 26.5 23.4 28.8 24.8
Table 3 presents classification error rates (severity level 5) on CIFAR-10C and CIFAR-100C. We group methods into two families: Test-Time Adaptation (TTA) and gradual domain adaptation (GDA). For TTA, “Source only” refers to the fixed pretrained model, and BN-1 updates batch normalization statistics on each test batch. The other baselines, including TENT-continual [43], AdaContrast [44], CoTTA [45], and GTTA-MIX [46], perform online adaptation of either the feature extractor or the classifier. Our approach, SWAT, combines the stability of batch-norm re-estimation with sample-wise alignment. Across both benchmarks, SWAT attains the lowest mean error (15.4% on CIFAR-10C, 24.8% on CIFAR-100C), outperforming the strongest prior TTA (GTTA-MIX: 15.6%/28.9%) and GDA competitors on nearly every corruption type.
5.3 Domain Shifts Analysis
[…]
00.20.20.40.40.60.60.80.81100.050.050.10.1distance to $\mathcal{H}_{0}$distance to $\mathcal{H}_{n}$distance to $\mathcal{H}_{l},\mathcal{H}_{r}$
Figure 4: The figure illustrates how the interpolated domains evolve in domain discrepancy along the adaptation path. The horizontal axis denotes the interpolation place $z\in[0,1]$ (0 = source domain, 1 = target domain). The vertical axis represents the $\mathcal{A}$-distance [47, 48], a proxy for distribution divergence. On the left, the $\mathcal{A}$-distance is computed with representations of a fixed encoder, while on the right, the distance is calculated using our SWAT representation.
Quantitative Analysis of Domain Shifts
We employ $\mathcal{A}$-distance [47] as the proxy of $\mathcal{H}\Delta\mathcal{H}$ distance to quantitatively evaluate the domain shifts, as illustrated in Fig. 4. We observe that the $\mathcal{A}$-distance between the source domain $H_{0}$ and the target domain $H_{n}$ exhibits large fluctuations (peak at 1.498), which indicates that directly aligning two domains causes unstable transfer or even negative transfer when the domain shifts are significantly large. In contrast, the proposed SWAT maintains near-zero distances (<0.11<0.11) to critical intermediate domains $\mathcal{H}_{l},\mathcal{H}_{r}$ across all positions, achieving a 63.7% reduction in the average $\mathcal{A}$-distance between $\mathcal{H}_{0}$ and $\mathcal{H}_{n}$ (0.104 vs. 1.284), demonstrating smooth knowledge transfer. The symmetrical reduction of bidirectional distances confirms balanced adaptation between forward and backward domain transitions.
(a) Source space $\mathcal{H}_{0}$
(b) Target space $\mathcal{H}_{n}$ w/o flow matching
(c) Target space $\mathcal{H}_{n}$ w/ flow matching
(d) $\mathcal{H}_{0}$$\mathcal{H}_{n}$ w/o adversarial training
(e) $\mathcal{H}_{0}$$\mathcal{H}_{g}$$\mathcal{H}_{n}$ w/ adversarial training
Figure 5: t-SNE [49] visualization of feature space geometry under different domain adaptation strategies through two complementary perspectives on Rotated MNIST (with 4 intermediate domains).
Visualization Analysis of Domain Shifts
The t-SNE visualizations (Fig. 5) reveal the geometric impact of different strategies: (1) Direct mapping to $\mathcal{H}_{n}$ without flow matching (Fig. 5(b)) causes catastrophic cluster overlap, as rigid alignment disrupts local semantic structures. (2) SWAT (Fig. 5(c)) maintains a high percentage of $\mathcal{H}_{0}$’s cluster purity through flow matching that preserves isometric relationships between neighboring domains $\mathcal{H}_{l}\leftrightarrow\mathcal{H}_{r}$. (3) The non-adversarial path $\mathcal{H}_{0}\rightarrow\mathcal{H}_{n}$ (Fig. 5(d)) exhibits discontinuous jumps (Hausdorff distance 4.72), while our adversarial flow $\mathcal{H}_{0}\rightarrow\mathcal{H}_{g}\rightarrow\mathcal{H}_{n}$ (Fig. 5(e)) reduces trajectory fragmentation by 75.6% (Hausdorff 1.15). This geometric perspective demonstrates the Semantic invariance and topological continuity of SWAT in the feature space.